The probability that brand A coffeemaker will brew bitter coffee is 0.3 and the probability that brand B coffeemaker will brew bitter coffee is 0.2. Two cups of coffee are produced; one by coffeemaker A and the other by coffeemaker B. Find the probability that:
(a) Both cups of coffee are bitter.
(b) Only the cup from brand A is bitter.
(c) Exactly one cup is bitter.
(d) Neither cups of coffee is bitter.
Please help. My brain is dead and this is my last question on the homework. Here's what I got, but I don't think it's correct:
(a) .2 x .3 = .06 or 6%
(b) .3 or 30%
(c) ?
(d) .7 x .8 = .56 or 56%

Answer by stanbon(75887)

The probability that brand A coffeemaker will brew bitter coffee is 0.3 and the probability that brand B coffeemaker will brew bitter coffee is 0.2. Two cups of coffee are produced; one by coffeemaker A and the other by coffeemaker B. Find the probability that:
(a) Both cups of coffee are bitter.
(b) Only the cup from brand A is bitter.
(c) Exactly one cup is bitter.
(d) Neither cups of coffee is bitter.
Please help. My brain is dead and this is my last question on the homework. Here's what I got, but I don't think it's correct:
(a) .2 x .3 = .06 or 6%
OK
--------
(b) .3 or 30%
No: 0.3*0.8 = 0.24 or 24%
---------------------------------
(c) ?
P(neither bitter) = 0.7*0.8 = 0.56
P(both bitter) = 0.2*0.3 = 0.06
P(one bitter) = 1 - (0.56+0.06) = 1-0.62 = 0.38
==================================
(d) .7 x .8 = .56 or 56%
OK
==============
Cheers,
Stan H.
==============